### v2

# Elisabethstraße 59/III, 8010 Graz, Austria

Abstract
Diplomatics, the analysis of medieval charters, is a major field of research in which paleography is applied. Annotating data, if performed by laymen, needs validation and correction by experts. In this paper, we propose an effective and efficient annotation approach for charter segmentation, essentially reducing it to object detection. This approach allows for a much more efficient use of the paleographer’s time and produces results that can compete and even outperform pixel-level segmentation in some use cases. Further experiments shed light on how to design a class ontology in order to make the best use of annotators’ time and effort. Exploiting the presence of calibration cards in the image, we further annotate the data with the physical length in pixels and train regression neural networks to predict it from image patches.
Keywords:
Diplomatics Paleography Object detection Resolution.1 Introduction
1.1 Diplomatics
Diplomatics is the study of diplomatic charters, medieval documents of a legal nature having a highly regular form. While diplomatics is mostly focused on the content of the charters, many visual aspects are subject of investigation.
Sigillography (the study of diplomatic seals), forgery detection, document identification, dating, etc. are all strongly dependent on analysing visual attributes of the documents making image analysis tools indispensable. Although diplomatics data has been part of larger datasets of historical document images, pure image analysis on charters has not been that common. Standing out among recent studies, Boroş et al. [1] demonstrates that Named Entity Recognition directly on images works better than with a two stage pipeline and the work by Leipert et al. [8] which addresses the problem of class imbalance in the context of image segmentation. In Fig.˜1 two typical charters, one verso and one recto, can be seen.
1.2 Monasterium.net
Our work is motivated by the need to analyse online archives, in particular monasterium.net, which constitutes the largest collection of continental European archives. It is focused on central Europe, containing 170 archives in total from Austria, Czech Republic, Estonia, Germany, Hungary, Italy, North Macedonia, Poland, Romania, Serbia, Slovenia, and Slovakia. The data in such archives is highly inhomogeneous in nature and many details about the acquisition process are unavailable and can only be guessed at. Thus, there is a need to perform automatic image analysis in order to obtain a proper overview of the quality of such archives.
1.3 Challenges
At the same time, in traditional methods of diplomatics, much of the scientist’s attention and effort is given to each specific charter. In the late medieval period, charters grow to the hundreds of thousands, much more than a scholar can embrace. This makes the need of automatic analysis assistance unavoidable. Among humanists such automation is known by the term distant reading, it is used to signify automatic text analysis instead of analysing the original documents; in contrast with the manual analysis of each particular document. this notion also extends to visual analysis known as distant seeing. Distant reading and seeing, as opposed to typical image analysis, is practically impossible to scrutinise experimentally and the extent to which such methodologies are falsifiable is a nuanced discussion. These epistemological challenges mandate a very good understanding of the underlying data and biases in the employed methods. From an engineering perspective, we considered that the most precious resource is the diplomatist’s time and most decisions were made by trying to minimize the scholar’s time needed for annotation and then trying to maximize the use of the spent time.
The main contributions of this paper are as follows:
1.
We introduce a highly efficient data annotation tool called FRAT.111Software available under an open-source licence at https://github.com/anguelos/frat and in pypi https://pypi.org/project/frat/
2.
We present an object detection diplomatics dataset and report experiments using it. 222The code has been forked from the original YOLOv5 repository, the specific scripts used for the experiments are in the bin folder https://github.com/anguelos/yolov5/tree/master/bin
3.
We present a dataset with manually annotated image resolution estimates and present a model that learns to infer them.333The code used to run the resolution regression experiments, including links to the datasets, is available in a github repository https://github.com/anguelos/resolution_regressor
2 Layout Analysis
2.1 1000 Charter Dataset
In order to make sure we gain a proper understanding of the data available in monasterium.net, we randomly chose 1000 charters from it. The heterogeneous nature of the dataset is illustrated by Fig.˜2, in which we plot the distribution of width, height, and surface in pixels and mega-pixels, respectively. Additionally, the number of diplomatic seals occurring in each charter image is given.
The dataset not only aims at understanding the quality and nature of the diplomatics data in monasterium.net but also at supporting the creation of pipelines that will offer distant viewing and distant reading. Specifically, we model the layout analysis problem as an object detection problem. As our goal is to maximize the efficiency of diplomatic scholars, we devised an annotation strategy that would require as little time as possible per charter. Diplomatic charters are considered highly formulaic [3], commonly containing a single dominant text-block as can be seen in Fig.˜1. Since the images do not seem to need any dewarping, we modeled all layout annotation tasks as defining two-point rectangles, i. e., a rectangle extending from the top left corner to the bottom right corner. Each rectangle belongs to one of 11 specific classes. While these classes are from the point of view of annotation flat and independent, they imply a hierarchical structure:
"No Class": Reserved, should never occur normally.
"Ignore": Used for possible detections that should be tolerated but not required when doing performance evaluation. Might make sense to exclude this class of objects when training a model, e. g., weights placed on charters in order to hold it in place for photography.
"Img:CalibrationCard": Calibration cards are usually the 21 cm Kodak ones but not always; they usually occur on both sides of the charter.
"Img:Seal": Many charters contain one or more seals. Seals are of particular interest for sigillographers, i.e. a small sub-community of paleographers focused on the study of seals.
"Img:WritableArea": This class represents the background material on which a charter was written. It consists typically of parchment but paper is also common in late medieval times. The following classes can only reasonably occur inside this class.
"Wr:OldText": This class represents the typically single text block of the charter. This rectangle should aim to contain the whole tenor, if possible only the tenor of the charter.
"Wr:OldNote": This refers to pieces of text put there during the charter’s creation/early life that are not part of the tenor.
"Wr:NewText": This refers to pieces of text added at later times, typically by archivists. From a legal point of view it should be considered irrelevant.
"Wr:NewOther": This refers to items added to the charter after its creation that are not purely text, e. g., stamps and seal imprints.
"WrO:Ornament": This refers to items put on the charter at creation time for principally aesthetic purpose.
"WrO:Fold": This class is meant to indicate a typical charter element, the plica. A fold that is not strictly a plica but is of significant size is still considered a fold.
As the class names imply: calibration cards, seals, and writable areas can be perceived as children of the root element “image”, and all other classes are children of the writable area. This structural interpretation is suggested but not enforced in any way.
2.2 Object Detection
[…]
The mean average precision (mAP) for all classes at Intersection over Union (IoU) obtained is , while at IoU it is . These numbers demonstrate very good results, for some classes, the results are even nearly flawless. The confusion matrix is shown in Fig.˜3. We can see that classes Img:Seal, Img:WriteableArea, and Img:CalibrationCard all score at precision or higher. These classes are the ones for which annotators only need little interpretation. Class "Wr:OldText" is also performing fairly well with an accuracy of , especially if one takes into account that of missed objects are detected as "Wr:OldNote", which indeed can easily be confused. This class is also quite important as it is the one that detects text blocks, which can subsequently be segmented into text lines so that they can be passed to a handwritten text recognition (HTR) process.
Fig.˜4 shows the precision recall curves of each class along with their mAP at IoU. It can be observed that the mAP scores are significantly higher than the precision values reported in the confusion matrix.
2.3 Annotations with FRAT
We introduce a lightweight annotation tool for document image annotation. The tool can annotate rectangles by just two-points with an arbitrary set of classes. Other than a class, each rectangle can also have an associated Unicode transcription as well as a comment. The tool is called Fast Rectangle Annotation Tool (FRAT). It is web-based and written in Python and Javascript and therefore runs on Linux and Mac OsX natively, and on Windows using Docker. FRAT is available in pypi and can be easily deployed with pip. In order to accommodate this very generic data model, the tool has its own JSON-based file format reflecting these fundamental assumptions and nothing more. The User Interface (UI) is aimed at being extremely lightweight and fast, to a certain extent at the expense of precision. It is inspired by gaming interfaces, where all buttons are accessible as single keys on the keyboard, see Fig.˜5. Set operations such as rectangle union and subtraction allow the annotator to figure out efficient use patterns that best suit the domain of data employed.
A user-overridable JSON configuration file defines all user preferences as well as the class labels to be employed. All that is needed in order to begin a new project is to define the class labels in the config file. Class labels are considered by FRAT to be flat although they might be perceived to form a hierarchy For example, if a class is textline and another class isword, FRAT will ignore the fact that a textline consists of words but simple heuristics could be made to realise this hierarchical relation, but this hierarchical relation could be realized through fairly straightforwrd post-processing heuristics, that would create textlines out of word bounding boxes. Custom exporting to PAGE-XML [12] for text-styled text-lines is demonstrated in the code. Since object detection fits best, the scenario FRAT was built around, generic exporting to the MS-COCO dataset format [9]. YOLOv5 [7] is also available but no transcriptions are exported.
From a user interface perspective, FRAT features two modes of image annotation: The first mode, “Label”, works like a typical generic object annotation tool, with an extra option of transcribing or commenting each annotated object. Defining an object is done by a single drag along its diagonal. Single-key shortcuts allow for editing, manipulating, and navigating through all defined objects. The second mode, “Transcribe”, is aimed at transcribing text and reviewing object classes. In this mode a list of all defined objects, their classes, their transcriptions, and their comments is displayed in a scrollable list. This mode is most effective when annotating transcriptions of textlines or words, since allowing to transcribe something out of its context might make the groundtruthing more realistic not allowing annotators to use high-level semantic cues from the nearby text. FRAT has also been used to annotate data including class and transcriptions in [13].
3 Resolution Prediction
Given that many images of diplomatic charters contain calibration cards, it is possible to infer an image resolution approximation. This is quite useful as it allows for quality analysis of large archives but would also allow to associate visual objects of the charter seal, plica, etc. with physical world sizes. As is often the case, the data we are dealing with has been acquired in a non-standardized way and any information about the acquisition process is considered irrecoverable. Under the assumption that during acquisition the charter, the calibration card, and the seals lie on the same plane and that local depth deviations are negligible, we can use the marked distances on the calibration cards to get an estimate for the width and height of any pixel that lies on that plane. Assuming a view angle , and assuming that calibration cards are predominantly on the horizontal edge of the cameras field of view, the estimated pixel resolution lies in between and of the estimated pixel resolution. In the case of a focal length, this error would be at most while in the case of focal length, it would be at most . As the images have been cropped, and nothing is known about this process, this distortion is not reversible but can be used to provide some bounds.
In order to create the ground truth data for the task of resolution estimation, we used the object detector to isolate the calibration cards, writeable areas, and seals. For simplicity’s sake, we used FRAT and marked rectangles with a diagonal wherever that was possible. When was not an option due to image size, we marked . In the few cases where there was no visible inch or cm marking, we operated on the assumption that inches are divided by halves, quarters, and eights; while centimeters are divided by halves and millimeters. For all analyses, we used Pixel per Centimeter (PpCm) unit.
3.1 Baseline
For a baseline, we considered the calibration-card object detector described in Section˜2. Specifically, since almost all calibration cards are variants of the Kodak calibration cards, they have a length of , even if they differ in other details, while having a height of less than . The calibration cards occur always on the sides of the images vertically or horizontally aligned either vertically or horizontally, by choosing the largest side of the bounding box, we can with high certainty obtain an estimate on how many pixels make a length of 20.5 cm on the photographed plane. This heuristic interpretation of the calibration card object detection gave a Spearman’s correlation of 89.13% and makes it a reasonably strong baseline. To the knowledge of the authors there is no other baseline available in the literature that is applicable in this experimental context.
3.2 ResResNet
We introduce a network called ResResNet which is a typical ResNet [4] where the final fully connected layer is substituted with a small multi-layer perceptron (MLP) consisting of three fully connected layers with 512, 256, and 1 nodes. While the ResResNet can utilise any ResNet architecture as a backend, we mostly experimented with ResNet18 in order to limit the capacity of the model to suit our relatively small dataset. Due to their average pooling layer, ResResNet can operate on images of arbitrary size. Image resolution can be considered as a texture-base analysis. Thus, during inference, we run ResResNet on a sliding window over the initial image and consider the median prediction as a result for the image. Note: in Fig.˜7 this inference method is called stable.
3.3 Training
[…]
3.4 Results and Analysis
Resolution estimation is a regression task. Hence, typical performance metrics such as Accuracy and F-score are not applicable. For our performance analysis, we use MSE and and we consider Spearmann’s correlation to be more appropriate as the errors do not follow a Gaussian distribution in a very loose sense. MSE quantifies the size of mistakes in total while Spearman’s correlation quantifies how often predictions are correct. Figure˜7 shows the results. While both validation and train sets were curated such that they would only refer to images that contained a calibration card out of necessity, they fail to demonstrate the base limitation of the baseline system, which is that a calibration card is required for an estimate to be possible at all.
Therefore, we gradually remove samples from the validation set from the most erroneous one (according to the squared error) to the least erroneous one. The results are shown in Fig.˜8. We notice that the calibration card baseline performs the best for most samples but overall ResResNet is more reliable as it handles hard samples more gracefully. Furthermore, ResResNet deals with images that do not contain any calibration card at all. The fact that “texture” analysis is enough to learn to predict physical resolution indicates that historical textual data is by nature scale-sensitive and because the trained image analysis models operate on various scales, we can speculate that they implicitly learn multiple scale representations of the same phenomenon. Further experiments that might help validate these hypotheses would be to see whether re-scaling train and test sets for typical textural image analysis tasks such as binarization would improve performance on these tasks. Another approach to better make use of this hypothesis would be to add ResResNet as a Spatial Transformer Network (STN) [5] with PpCm being the single degree of freedom.
4 Discussion and Conclusion
4.1 Data Annotation Efficiency
In this paper, we present a work on Diplomatic charters motivated by making optimal use and re-use of domain experts. Our principal research question was how much effort can we economise on annotating and how far will these weak annotations take us in the use of subsequent pipelines. While observing the annotation task, it was evident that marking the seals on a charter would only take a few seconds while waiting for the image to load might take quite a bit longer. To a great extent the proposed class ontology was designed to make better use of the scholar’s time by trying opportunistically to ask more questions while they wait. Essentially we tried to increase the objects to be annotated on each image in order to dilute waiting time of annotators to the extent where it would not cause frustration.
The defined classes were chosen not only on the theoretical completeness but also on what is easy to obtain and what might be easier to learn from a machine learning perspective. Modeling the classes and the annotation recommendation benefited greatly from the interdisciplinary nature of the team. It is worth pointing out that the class that made diplomatists sceptical was the class Wr:Fold. Since it does not coincide totally with the established “plica” charter element. It is also the worst performing one, while this could also be interpreted with the fact that folds are not well defined in rectangles and therefore object detection might possibly not be well suited for these elements.
4.2 Towards a Full Diplomatics Pipeline
It also appears that in the case of diplomatics, there is no need for a dewarping stage as in other DIA (Document Image Analysis) pipelines [10]. While dewarping is a very simple step on its own, it complicates a DIA pipeline quite a bit because pixel coordinates are not consistent across the whole pipeline and from a user interface perspective it becomes a major challenge to align localised outputs to the original images. The performance trade-off of using simple object detection as the only supervision data for layout analysis in other use cases has yet to be determined. Thus, there is strong evidence that it is sufficient in the diplomatics case.
4.3 Future Work
[…]
Acknowledgements
The work presented in this paper has been supported by ERC Advanced Grant project (101019327) "From Digital to Distant Diplomatics" and by the DFG grant No. CH 2080/2-1 "Font Group Recognition for Improved OCR".